OpenStudy (anonymous):

Write the equation of the line that passes through (-4,12) and (7,23)

OpenStudy (anonymous):

Those are the choices i have A. y = x + 16 B. y = 3x + 18 C. y = 1/6 x + 16 D. y = 1/3 x + 18

OpenStudy (anonymous):

Find the slope first, \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] m is the slope. Next you want yo put it in y = mx+b form, you can do it by looking at it but you can also use point - slope form, do some algebra and put it into that form. \[y-y _{1}=m(x-x _{1})\] - point slope form Your \[(x_1,y_1)\] here are your first coordinates (-4,12) You should be able to do it from here, and if you would like more help just tag me, but make sure you show some work before you want further help. Good luck.